OpenStudy (anonymous):

sound intensity of noise is 100dB.find sound intensity sound intensity of noise is 100dB.find sound intensity @Physics

OpenStudy (anonymous):

10log(I/Io)=db and Izero=Io=10^(-12) W/m^2 base of the log is 10 just put the numbers in 10(logx/10^(-12))=100db after the simple log algebra ull get the x